OpenStudy (anonymous):

what are the roots of the equation −x2 − 2x + 8 = 0? How do you solve this?

OpenStudy (angela210793):

D=b^2-4ac=4-4*(-1)*8=36>0 x1=(-b+sqrD)/2a=2+6/-2=-4 x2=(-b-sqrD)/2a=(2-6)/-2=2

OpenStudy (amistre64):

it takes a good eye, or a remembrance of the quadratic formula

OpenStudy (anonymous):

it is \[-(x ^{2}+2x-8)=-(x ^{2}+4x-2x-8)=-(x+4)(x-2)\] hope its correct

OpenStudy (anonymous):

roots are -4,2

OpenStudy (anonymous):

how'd you get the 4x